OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)(2x2y6)

OpenStudy (mertsj):

\[(3x^3y^4)(2x^2y^6)=(3\times 2)(x^3\times x^2)(y^4\times y^6)\]

OpenStudy (mertsj):

Think of it like that.

OpenStudy (anonymous):

Oh alright

OpenStudy (anonymous):

so

OpenStudy (anonymous):

6x5y10?

OpenStudy (mertsj):

Good job!!

OpenStudy (anonymous):

Thanks :) Ur awesome!!!!!

OpenStudy (anonymous):

I got a few more questions after this if u could help me....